OpenStudy (anonymous):

Solve using substitution: 2x = 1 - y 5x - 10y = -10 HELP!!! Will medal and fan for solution

OpenStudy (anonymous):

So I find it easier to flip them so they are facing the same. (Meaning have y= on one side and the rest on the other.) 2x + y = 1 5x - 10y = -10 Since they dont have an opposite quality, i need to multiply one equation to make it to where i can cancel out a number. 2x + y = 1 <-- (*10) 5x - 10y = -10 This would turn it into: 20x + 10y = 10 5x - 10y = -10 Cancel out the 10y and -10y to get: 20x = 10 5x = -10 Add them 25x = 0 x = 0 Then plug into an original equation and find y: 2x + y = 1 2(0) + y = 1 0+y=1 y = 1 and x = 0 (0,1) is your answer

OpenStudy (anonymous):

I hope that makes sense. All you have to do is make the x values or y values with a positive in one equation and a negative in the other. Then just completely cancel them both and combine the equations.

OpenStudy (anonymous):

Yes it does kind of...Thank you!!:) Now how do I medal you? ha

OpenStudy (anonymous):

Never mind I figured it out! Thanks again

OpenStudy (anonymous):

No problem! These are my favorite kind of questions so message me if you need help again!